OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (1, -1) and a directrix y = 4.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

ha..

OpenStudy (anonymous):

Im ur first fan

OpenStudy (anonymous):

y=1=-1/20(x-1)^2 y+1=1/20(x-1)^2 y+1=-1/12(x-1)^2 x+1=-1/20(y-1)^2

OpenStudy (anonymous):

your def not my first..but thanks

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

hmm - can't remember this stuff - have to look it up

OpenStudy (anonymous):

no im ur first fan coolaidd. ohhhh yeaahh

OpenStudy (cwrw238):

nope - sorry - its been too long a time coolaid

OpenStudy (anonymous):

ughh..its ok i guess